OpenStudy (annie96):

Can anyone tell me that i got the answers right. 3. Solve the equation for y and identify the slope: 2x + y = 5 (Points : 1) ( 2) Answer -2 1/2 5 Question 4.4. Solve the equation for y and identify the slope: -2x + 2y = 4 (Points : 1) -2 1 (-1) Answer 2 Question 5.5. Solve the equation for y and identify the slope: x + 2y = 6 (Points : 1) 1 ( 2) Answer -1/2 -1

jimthompson5910 (jim_thompson5910):

All 3 are incorrect

OpenStudy (annie96):

Thank you.

jimthompson5910 (jim_thompson5910):

The idea is to solve for y and get it into the form y = mx+b. So let's say we want the slope for x + 2y = 6 x + 2y = 6 2y = 6 - x 2y = -x + 6 y = (-x + 6)/2 y = (-x)/2 + 6/2 y = (-1/2)x + 3 This is now in the form y = mx+b where m = -1/2 and b = 3 m = -1/2 is the slope b = 3 is the y-intercept, which is the point (0,3)

OpenStudy (annie96):

ok thank u

jimthompson5910 (jim_thompson5910):

you're welcome